perform matrix row operation and write the new matrix
1 2 2 2
0 1 -1 2 -5R2+R3
0 5 4 1

Answer by Edwin McCravy(20056)

perform matrix row operation and write the new matrix
1  2  2  2
0  1 -1  2   -5R2+R3
0  5  4  1

-5R2+R3
 
-5R2 means to do this work to the side 

1. Multiply row #2 by -5 :  -5(0  1 -1  2) = (0 -5  5 -10)

and then +R3 means
 
2. Add that to row #3  :     (0 -5  5 -10) + (0  5  4  1) = (0  0  9  -9)
         
3. Replace row #3 by (0  0  9  -9)

1  2  2  2
0  1 -1  2   
0  0  9 -9

That's the new matrix.

Edwin